OpenStudy (anonymous):

Line integrals. I have to evaluate a line integral along the curve of a parabole x=t, y=t²+1. I'm not sure how to parametrize this.

OpenStudy (anonymous):

The integral given is \[\int\limits_{(01)}^{(1,2)}[(x^2-y)dx+(y^2+x)dy]\]

OpenStudy (anonymous):

Do I just plug in the t values for the x and y as well as the derivatives?

OpenStudy (rational):

yes, x = t dx = ? y = t^2+1 dy = ?

OpenStudy (anonymous):

sorry that lower bound ought to be (0,1) dx=1 dy=2t

OpenStudy (rational):

dx=1 dt dy=2t dt

OpenStudy (anonymous):

but don't I have to parametrize the bounds?

OpenStudy (rational):

\[\large \int\limits_{(0,1)}^{(1,2)}[(x^2-y)dx+(y^2+x)dy]\] after parameterization : \[\large \int\limits_{0}^{1}[(t^2-(t^2+1))dt+((t^2+1)^2+t)2tdt]\]

OpenStudy (anonymous):

how come it's \[0 \le t \le 1\]

OpenStudy (kainui):

Because your vector is really: \[\LARGE \bar v(t)=<x(t), y(t)>=<t, t^2+1>\] So now let's evaluate v(t) at the end points what do we get? \[\LARGE \bar v(0)=<0, 1> \\ \LARGE v(1)=<1,2>\] Similarly if we just graphed the non parameterized parabola y=x^2+1 we would get the same ordered pairs! Any time you want to parameterize a function you can simply just put in x=t and replace all the x's in the function y(x) with t's to get y(t). Play around with this a little bit and I think it will become clear how you could turn any function into a function of a single parameter. From a vector point of view, all along you were really just looking at a function where x only scaled by 1 every time, hence x(t)=t!

OpenStudy (anonymous):

Thanx, @kainui, and @rational. I found 35/3.